# Lower tensor-to-scalar ratio as possible signature of modified gravity

### v1

@@ III Embedding as a modified gravity model
The use of the set of perturbed equations in Eqs.(55), (56, (57) and (58) is simplified after some algebra into the following set of equations
@@
where $\rho\Delta=\bar{\rho}\delta+3\frac{\mathcal{H}}{k}(\bar{\rho}+P)\theta$. The set of Eqs.(60) and (61) are valid for all times. Thus, $\mu(a,k)$ and $\Sigma(a,k)$ functions can be written as
$\mu(a,k)=\frac{1}{1-\frac{\chi(a)}{k^{2}}}\left[1+\frac{3(\bar{\rho}+P)\sigma}{\rho\Delta}\right]\;,$ (62)
@@
It is worth noting that the anisotropic stress is inserted in $\mu$ function. This is due to the fact that we do not define an extrinsic source, rather, the influence of extrinsic geometry is done over the matter content perturbing background geometry. Thus, the anisotropy, if considered, is a full matter-source.
@@
When anisotropic stresses are neglected, one does not obtain the standard GR limit as $\Sigma(a,k)=\mu(a,k)$ and $\gamma(a,k)=1$. In this framework, the equality of Newtonian potentials $\Phi=\Psi$ is obtained in twofold conditions: anisotropic stress must vanish and the extrinsic term $\gamma_{s}\rightarrow 0$ in order to recover GR correspondence.
The set of basic equations are complete with the matter perturbation equations in Eqs.(39) and (39), and the evolution equation of $\Phi$. To obey solar constraints $\gamma_{s}$ in $\mu(a,k)$ function must comply with the condition
@@ IV Extrinsic curvature as an effective inflaton field
with $\kappa=8\pi G$ and $\bar{\rho}=\bar{\rho}_{m}+\bar{\rho}_{rad}$. In terms of inflationary cosmology, such energy density $\bar{\rho}_{ext}$ should provide a response in a form of a scalar field potential energy density $V(\phi)$ generated by a spatially homogeneous “extrinsic” scalar field $\phi$ [50]. Thus, one defines a Lagrangian $\mathcal{L}_{\phi}$ such as
@@
where $\bar{\rho}_{ext}$ and $\bar{p}_{ext}$ denote the energy density and pressure generated by extrinsic curvature as a function of the extrinsic scalar field $\phi$.
@@
where $\alpha_{0}=\sqrt{|3(1+w)|}$. This kind of exponential model is well-known in literature commonly referred as power law inflation (PLI)[65, 66, 67, 68]. Such potential was also studied in the context of M-theory [69] and RSII scenarios [70]. Due to the strong constraints imposed by Planck data [71] on PLI, we follow the generalization proposed in Refs.([72, 73]) named $\beta$-exponential inflation. In Ref.([72]), the authors introduced a class of potentials in a form
@@ V Data and methodology
To delineate constraints on the free parameter of the theory framework under consideration, we will examine several datasets individually as well as in various combinations. Specifically, we will consider:
  • Utilizing the latest NPIPE Planck DR4 likelihoods [54, 55] of Planck 2018 legacy data release, our analysis incorporates Cosmic Microwave Background (CMB) measurements. These include the $-\ell$ Plik TT likelihood spanning the multipole range $30\leq\ell\leq 2508$, as well as TE and EE measurements within the multipole range $30\leq\ell\leq 1996$. Additionally, we incorporate $-\ell$ TT-only ($2\leq\ell\leq 29$) and EE-only ($2\leq\ell\leq 29$) likelihoods. Furthermore, our dataset encompasses CMB NPIPE Planck lensing power spectrum measurements, collectively referred to as the Planck dataset.
@@
    $f(z)=\frac{\partial\ln\sigma_{8}}{\partial\ln a}.$ (90)
    The RSD measurements provide constraints on the quantity $f(z)\sigma_{8}(z)$. The $\sigma_{8}(z)$ depends on the matter power spectrum, $P(k,z)$, which is calculated by default in the Boltzmann code. Both $f(z)$ and $\sigma_{8}(z)$ are sensitive to variations in the effective gravitational coupling and the light deflection parameter, which play a crucial role in our Poisson and lensing equations in MG models.
[…]
Parameter Planck+BK18 Planck+BK18+LSS
[…]
$r$ $0.0142^{+0.0049}_{-0.012}$ $<0.0307$ $0.0139^{+0.0044}_{-0.013}$ $<0.0303$
@@
[…]
@@ VI Results
We commence our analysis of the results by examining the parameters that indicate deviations from the standard $\Lambda$CDM cosmology. Figure (2) illustrates the parameter space for $\gamma_{s}$, $\mu$, and $\Sigma$.
@@
[…]
@@ VII Final Remarks
In this paper, we have derived the gravitational equations within four dimensions by inducing them from a five-dimensional bulk, employing Nash’s embeddings framework and incorporating them into a well-established $\mu$ $\Sigma$ MG representation. From the analysis of the slow-roll conditions, we have obtained a PLI model that was generalized to the $\beta$-exponential inflation [72, 73] that suggest $w=-0.9944$ for the number of e-folds in the range $N=50$ and $N=60$. Interestingly, this model also sits within the previous reasonable expected number e-folds compatible with the tighter restriction on $r$, which is important to maintain the window of solving the horizon problem. This enables us to assess and investigate the impact of the model on linear perturbations in the CMB data. Apart from the values of the tensor-to-ratio parameter, our primary analyses revealed similar predictions to the ones of the $\Lambda$CDM model. Additionally, we have quantified the model’s predictions concerning inflationary dynamics. By utilizing data from CMB-PR4, BICEP/Keck Array 2018, and certain LSS measurements, we have established an tighter upper limit of $r<0.0303$ at a 95% C.L. As compared with $\Lambda$CDM model, such apparent improvement of tensor-to-ratio parameter and a similar but lower value of the scalar spectral index $n_{s}$ values may suggest a glimpse of a modified gravity signature which may be improved in future experiments.

### v2

@@ III Embedding as a modified gravity model
The use of the set of perturbed equations in Eqs.(63), (64, (65) and (66) is simplified after some algebra into the following set of equations
@@
where $\rho\Delta=\bar{\rho}\delta+3\frac{\mathcal{H}}{k}(\bar{\rho}+P)\theta$. The set of Eqs.(68) and (69) are valid for all times. When anisotropic stress is neglected, $\mu(a,k)$ and $\Sigma(a,k)$ functions can be written as
$\mu(a,k)=\frac{1}{1-\frac{\chi(a)}{k^{2}}}\;,$ (70)
@@
When the extrinsic term $\gamma_{s}\rightarrow 0$ in order to recover GR correspondence, one obtains the standard GR limit as $\Sigma(a,k)=\mu(a,k)$ and $\gamma(a,k)=1$. Thus, the set of basic equations is complete with the matter perturbation equations in Eqs.(47) and (47), and the evolution equation of $\Phi$. To obey solar constraints, $\gamma_{s}$ in $\mu(a,k)$ function must comply with the condition
@@ IV Extrinsic curvature as an effective inflaton field
with $\kappa=8\pi G$ and $\bar{\rho}=\bar{\rho}_{m}+\bar{\rho}_{rad}$. In terms of inflationary cosmology, such energy density $\bar{\rho}_{ext}$ should provide a response in a form of a scalar field potential $V(\phi)$ generated by a spatially homogeneous “extrinsic” scalar field $\phi$ [50]. Thus, one defines a Lagrangian $\mathcal{L}_{\phi}$ such as
@@
where $\bar{\rho}_{ext}$ and $\bar{p}_{ext}$ denote the energy density and pressure generated by extrinsic curvature as a function of the extrinsic scalar field $\phi$, respectively.
@@
where $\alpha_{0}=\sqrt{|3(1+w)|}$. This kind of exponential model is well-known in literature and is commonly referred as power law inflation (PLI)[70, 71, 72, 73]. Such potential was also studied in the context of M-theory [74] and RSII scenarios [75]. Due to the strong constraints imposed by Planck data [76] on PLI, we follow the generalization proposed in Refs.([77, 78]) named as $\beta$-exponential inflation. In Ref.([77]), the authors introduced a class of potentials in a form
@@ V Data and methodology
Table 2: Summary of $68\%$ and $95\%$ confidence level limits for the parameters of interest obtained from Planck+BK18 and Planck+BK18+LSS joint data at pivot scale $k_{p}=0.05$Mpc-1.
Parameter Planck+BK18 Planck+BK18+LSS
[…]
$r$ $0.0142^{+0.0049}_{-0.012}$ $<0.0307$ $0.0139^{+0.0044}_{-0.013}$ $<0.0303$
To delineate constraints on the free parameter of our model, we examine several datasets individually as well as in various combinations. Specifically, we will consider:
  • Utilizing the latest NPIPE Planck DR4 likelihoods [54, 55] of Planck 2018 legacy data release, our analysis incorporates Cosmic Microwave Background (CMB) measurements. These include the $-\ell$ Plik TT likelihood spanning the multipole range $30\leq\ell\leq 2508$, as well as TE and EE measurements within the multipole range $30\leq\ell\leq 1996$. Additionally, we incorporate $-\ell$ TT-only ($2\leq\ell\leq 29$) and EE-only ($2\leq\ell\leq 29$) likelihoods. Furthermore, our dataset encompasses CMB NPIPE Planck lensing power spectrum measurements, collectively referred as the Planck dataset.
@@
    $f(z)=\frac{\partial\ln\sigma_{8}}{\partial\ln a}.$ (98)
    The RSD measurements provide constraints on the quantity $f(z)\sigma_{8}(z)$. The $\sigma_{8}(z)$ depends on the matter power spectrum, $P(k,z)$, which is calculated by default in the Boltzmann code. Both $f(z)$ and $\sigma_{8}(z)$ are sensitive to variations in the effective gravitational coupling and the light deflection parameter, which play a crucial role in Poisson and lensing equations in MG models.
@@
Here, $\Omega_{DE}(a)$ is denoted as the extrinsic contribution $\Omega_{ext}(a)$ which means that $\Omega_{DE(0)}=\Omega_{ext}(0)=1-\Omega_{m(0)}$. The form of Eqs. (99) and (100) in this parametrization is obtained from expanding the function $\chi(a)<<1$ in the denominator of Eq. (70). We assume a pivot scale fixed at $a_{p}=10^{-4}$ and $k_{p}=0.05$Mpc-1. It is worth noting that the anisotropic stress is not considered in our analysis.
@@ VI Results
We analyze the results by examining the parameters that indicate deviations from the stan standard $\Lambda$CDM cosmology. Figure (2) illustrates the parameter space for $\gamma_{s}$, $\mu$, and $\Sigma$.
@@
Figure (3) shows the $n_{s}-r$ plane. It is noteworthy to observe that the tensor-to-scalar ratio effects arise from the modified gravity scenario outlined in this study. In canonical parametric $P(k)$ inflation and $\Lambda$CDM dynamics, $r<0.036$ at 95% CL  [29] is found, whereas for the present modified scenario, we find a $r<0.0307$ at 95% CL with Planck+BK18 data. This represents a variation of $\Delta r=0.03$ from standard model tightening the constraint on $r$ suggesting a signature of a MG model rather than $\Lambda$CDM. At pivot scale $k_{p}=0.05$Mpc-1, our result is compatible with the upper bounds predicted by current Planck data [76] with $r<0.032$ at $95\%$CL from $E$ and $B-$mode spectra BK18 [56] and LSS data [79]. In addition, our upper constraint on $r$ with $r<0.0307$ at $95\%$CL is also tighter than the standard joint data Planck+BK18 with $r<0.036$ and NPIPE (PR4) with $r<0.056$ at 95% CL. Our results indicate a lower value for $r$ as compared with the observations from the PR4 BB spectrum for multipoles between $l=2$ and $l=150$ with $r=0.033\pm 0.069$ [76]. In contrast with the frequentist profile likelihood method [30] with an upper limit of $r<0.037$ at $95\%$ CL, we also obtain a tighter constraint on $r$ for the same combination of Planck, BK18 and LSS with $r<0.0303$ at $95\%$ CL. In this context, our inflationary model partially constraints data better than the $\Lambda$CDM model does.
@@ VII Final Remarks
In this paper, we have derived the gravitational equations within four dimensions by inducing them from a five-dimensional bulk, employing Nash’s embeddings framework and incorporating them into a well-established $\mu$-$\Sigma$ representation. From the analysis of the slow-roll conditions, we have obtained a PLI model that was generalized to the $\beta$-exponential inflation [77, 78]. We have obtained $w=-0.9944$ for the number of e-folds in the range $N=50$ and $N=60$. Interestingly, this model also sits within the previous reasonable expected number e-folds compatible with the tighter restriction on $r$, which is important to maintain the window of solving the horizon problem. This enables us to assess and investigate the impact of the model on linear perturbations in the CMB data. Apart from the values of the tensor-to-ratio parameter, our primary analyses revealed similar predictions to the ones of the $\Lambda$CDM model. Additionally, we have quantified the model’s predictions concerning inflationary dynamics. By utilizing data from CMB-PR4, BICEP/Keck Array 2018, and certain LSS measurements, we have established a tighter upper limit of $r<0.0303$ at 95% C.L. As compared with $\Lambda$CDM model, such apparent improvement of tensor-to-ratio parameter and a similar but lower value of the scalar spectral index $n_{s}$ values may suggest a glimpse of a modified gravity signature which may be improved in future experiments.